OpenStudy (anonymous):

(x^3)^5 A.x^8 B.x^15 C.x^2 D.5x^3

OpenStudy (ns36):

A. When you have a power question like this, you can simply add the powers, so because 3+5 = 8, the answer is x^8.

OpenStudy (anonymous):

when u have questions of this type multiply the powers. (a^b)^c=a^(bc) a^b*a^c=a^(b+c) a^b/a^c=a^(b-c) Hope that helps